74% of freshmen entering public high schools in 2006 graduated with their class in 2010. A random sample of 81 freshmen is selected. Find the probability that the proportion of students who graduated is greater than 0.743 .
Answer #1

Given,

p = 0.74 , n = 81

Using central limit theorem,

P( < p) = P( Z < - p / sqrt( p( 1 -p ) / n))

So,

P( > 0.743) = P( Z > 0.743 - 0.74 / sqrt( 0.74 * 0.26 / 81) )

= P( Z > 0.0616)

= 0.4755